OpenStudy (anonymous):

Establish the Identity: cot^2 (θ/2) = (secθ +1)/(secθ -1)

OpenStudy (anonymous):

(secθ +1)/(secθ -1) = (1/cosθ +1)/(1/cosθ -1) = (1+cosθ)/(1-cos θ) = (2cos^2θ/2 )/(2sin^2θ/2) = cot^2θ/2 [proved]